OpenStudy (anonymous):

3/4+7/8 its not necessary to simplify answer. please help i dont get it and explain

OpenStudy (anonymous):

\[\frac{3}{4}+\frac{7}{8}=\frac{2*3+7}{8}=\frac{13}{8}\]